Q1 What is the magnitude of the momentum of a fly if it is traveling with...

Q1

What is the magnitude of the momentum of a fly if it is traveling with speed 1 m/s and has a mass of 0.0001 kg?

Select one:

a. 0.0001 kg·m/s.

b. 1.0001 kg·m/s.

c. 6.0001 kg·m/s.

d. 5.0001 kg·m/s.

e. None of the above

Q2

A bullet fired from the muzzle of a rifle initially rotates with an angular velocity of 2200 radians/sec. If the bullet is slowed by friction and rotates through 480 revolutions during its 1.4 seconds of flight, what is the angular acceleration of the bullet?

Select one:

a. 0.98 rad/s2

b. 960ð rad/s2

c. 65.38 rad/s2

d. 25.38 rad/s2

e. -65.38 rad/s2

Q3

Four objects are situated along the y-axis as follows: a 2 kg object is at +3 m, a 3 kg object is at +2.5 m, a 2.5 kg object is at the origin, and a 4 kg object is at -0.5 m. Where is the center of mass of these objects?

Select one:

a. (xCM, Ycm)= (1,0)

b. (xCM, Ycm)= (0,1)

c. None of the above

d. (xCM, Ycm)= (0,0)

e. (xCM, Ycm)= (1,1)
Answer #1

Use the definition of momentum and equation for rotational kinematics to find the solution as shown belowP = mv = (0-0001 kg)mls) = 10.0001 19m/s o = Witt 1 2+² => 480x27 rad = (2200 rad)){-45) + Ź«(2:45) ? + -65.38) vad/s2 3) sin

***********************************************************************************************
This concludes the answers. If there is any mistake or omission, let me know immediately and I will fix it....